# Improving noisy free-energy measurements by adding more noise

### v2

@@ Abstract
Estimating free-energy differences using nonequilibrium work relations, such as the Jarzynski equality, is hindered by poor convergence when work fluctuations are large. For systems governed by overdamped Langevin dynamics, we propose the counterintuitive approach of adding noise in order to increase the precision of such calculations. By introducing additional stochastic fluctuations to the system and rescaling its potential energy, we leave the thermodynamics of the system unchanged while increasing its relaxation rate. For a given time-dependent protocol this modification reduces dissipated work, leading to more accurate free-energy estimates. We demonstrate this principle using computer simulations applied to two model systems. However, the regime of applicability of this strategy is likely limited, because it requires control of the system’s potential energy in a way that is feasible in only a few experimental settings.
@@ I Introduction
The Jarzynski equality and the Crooks fluctuation theorem provide a way to estimate equilibrium free-energy differences from nonequilibrium work measurements Jarzynski (1997); Crooks (1999). These relations apply to systems initially at equilibrium and driven out of it by the variation of a set of external control parameters. However, such estimates can be numerically challenging. The Jarzynski equality involves an exponential average, and rare trajectories with atypically small work values contribute disproportionately, making free-energy estimates unreliable unless a large number of trajectories are sampled Jarzynski (2006); Hummer and Szabo (2010); Rohwer et al. (2015). This problem worsens as the rate of driving increases, because the mean dissipated work $\langle W\rangle-\Delta F$ grows, leading to an exponential increase in the number of trajectories required to estimate the free-energy difference to a given precision. Various strategies have been explored to mitigate this issue, including optimizing control protocols to minimize dissipated work and work fluctuations Schmiedl and Seifert (2007); Geiger and Dellago (2010); Vaikuntanathan and Jarzynski (2008); Blaber and Sivak (2020); Arrar et al. (2019); Loos et al. (2024); Zhong and DeWeese (2022).
@@
In this paper we propose an alternative and counterintuitive approach to increasing the precision of free-energy measurements for systems governed by overdamped Langevin dynamics: add noise. By injecting additional stochastic fluctuations into the system and rescaling its potential energy, we can increase the system’s relaxation rate without altering its thermodynamics. For fixed driving rate, this modification reduces dissipative work and so suppresses rare-event sampling problems, leading to more precise estimates of $\Delta F$. We demonstrate this effect in two model systems using numerical simulations, showing that this noise-engineering strategy can significantly enhance the accuracy of free-energy calculations. Our results suggest that noise engineering could be a valuable tool for enhancing thermodynamic measurements and computations, and that the notion of optimal control could be broadened to include control of noise, which is now experimentally feasible Mart nez et al. (2012); Chupeau et al. (2018); Saha et al. (2023); Aifer et al. (2024), as well as control of the potential.
@@
However, the regime of applicability of this method to experimental free-energy calculations is limited. Noise-injection methods have been demonstrated in various experimental settings, but the method discussed here also depends on our ability to control the system’s potential energy, which is feasible in only a few experimental settings. We describe the method in Section II, and apply it to two model systems in Section III. We conclude in Section IV, where we comment on the applicability of the approach to various experimental systems.
@@ II Free energies from nonequilibrium measurements
$\dot{x}_{i}=-\mu\frac{\partial}{\partial x_{i}}U({\bm{x}},{\bm{c}}(t))+\sqrt{2k_{\rm B}T\mu}\,\eta(t),$ (1)
on the time interval $t\in[0,t_{\rm f}]$. Here $\mu$ is the mobility parameter, which sets the basic time scale of the system, and $\eta$ is a Gaussian white noise with mean $\langle\eta(t)\rangle=0$ and covariance $\langle\eta(t)\eta(t^{\prime})\rangle=\delta(t-t^{\prime})$. The noise $\eta$ represents the thermal fluctuations inherent to the system.
@@
[…]
@@
where $i$ labels trajectories, as well as a measure of the statistical error in this quantity, the variance $\sigma^{2}_{B}$ of the block average
$B=N_{\rm block}^{-1}\sum_{i=1}^{N_{\rm block}}{\rm e}^{-\beta W_{i}}.$ (7)
We can also calculate the number of trajectories required to estimate $\Delta F$ to a particular precision, say $0.1\,k_{\rm B}T$, using the formula Geiger and Dellago (2010)
$N_{0.1}=100\left(\langle{\rm e}^{-2\beta W}\rangle-1\right).$ (8)
@@
$\dot{x}_{i}=-\mu\frac{\partial}{\partial x_{i}}\lambda U({\bm{x}},{\bm{c}}(t))+\sqrt{2k_{\rm B}T\mu}\,\eta(t)+\sigma\zeta(t),$ (9)
where $\langle\zeta(t)\rangle=0$ and $\langle\zeta(t)\zeta(t^{\prime})\rangle=\delta(t-t^{\prime})$. The two noise terms in (10) can be considered an effective Gaussian white noise of variance $2k_{\rm B}T\mu\lambda$, and so (9) can also be written
$\dot{x}_{i}=-\mu\frac{\partial}{\partial x_{i}}U_{\lambda}({\bm{x}},{\bm{c}}(t))+\sqrt{2k_{\rm B}T_{\lambda}\mu}\,\eta(t),$ (10)
@@
[…]
@@
This effective time-rescaled description suggests that the $\lambda$-modified dynamics will allow more accurate estimation of $\beta\Delta F$ using the Jarzynski identity than does the original dynamics. From our understanding of how work scales with trajectory duration Schmiedl and Seifert (2007); Van Vu and Hasegawa (2021) we expect the dissipated work $\beta_{\lambda}\langle W_{\lambda}\rangle_{\lambda}-\beta\Delta F$ to decrease with $\lambda$, and for large $\lambda$ to scale as $1/\lambda$. Reducing the dissipated work tends to reduce the variance of the Jarzynski estimator, and so we expect the statistics of $\langle{\rm e}^{-\beta_{\lambda}W_{\lambda}}\rangle_{\lambda}$ to become increasingly better behaved as $\lambda$ increases.
@@ III.1 Particle in a trap
Figure 2: Trap-translation model using the linear protocol within the $\lambda$-modified dynamics (9) (green lines). The trajectory duration is $t_{\rm f}=1$, as in Fig. 1. Dashed lines labeled “opt” correspond to the optimal protocol used within the original dynamics (for which $\lambda=1$). (a) Mean work $\langle W_{\lambda}\rangle_{\lambda}$, in units of $k_{\rm B}T$, as a function of $\lambda$. (a) Mean work $\beta_{\lambda}\langle W_{\lambda}\rangle_{\lambda}$ as a function of $\lambda$. As expected from the arguments connecting Eq. (9) and Eq. (14), this quantity behaves as $\beta\langle W\rangle$ for the original dynamics, Eq. (1), if the protocol is imposed $\lambda$ times more slowly for $\lambda$ times longer. The function $w$ is given by Eq. (17). (c) Free-energy estimator (13). (d) Variance of the block average (7), using $N_{\rm block}=100$. (e) Number of trajectories (8) required to estimate $\Delta F$ to a precision $0.1k_{\rm B}T$. (f) Work distributions for forward and reverse protocols, for $\lambda=30$; compare the original dynamics of Fig. 1(b). Averages and distributions are calculated using $N_{\rm traj}=10^{6}$ trajectories.
@@
Figure 5: Erasure using the basic protocol of Ref. Dago et al. (2023) within the $\lambda$-modified dynamics (9). (a) Time-dependent protocol ${\bm{c}}(t)$. (b) Probability of successful erasure as a function of $\lambda$. (c) Free-energy estimator (13). (d) Variance of the block average (7). Averages are calculated using $10^{5}$ independent trajectories.
@@
For this system the optimal protocol is also the protocol that minimizes the error in estimating (3), and it possesses a Gaussian work distribution Geiger and Dellago (2010).
@@
which is slightly larger than that of the optimal protocol for all $t_{\rm f}$ Schmiedl and Seifert (2007). For large $t_{\rm f}$, both $\beta\langle W\rangle_{\rm opt}$ and $\beta\langle W\rangle_{\rm lin}$ go to zero as $\sim kc_{\rm f}^{2}/t_{\rm f}$.
@@
In Fig. 2 we show results obtained using the $\lambda$-modified dynamics (green lines), together with those obtained using the optimal protocol. Panel (a) shows that the mean work $\langle W_{\lambda}\rangle_{\lambda}$ increases with $\lambda$, as might be expected from the use of an energy function $\lambda$ times larger than the original. The new dynamics is more costly energetically than the original dynamics, even without accounting for the energy required to inject the added noise.
However, consistent with our expectation from the previous section, the combination $\beta_{\lambda}\langle W_{\lambda}\rangle_{\lambda}$ decreases with increasing $\lambda$, because the system behaves as if it is subject to a slowed-down version of the original protocol. This decrease is shown in Fig. 2(b). Indeed, we see that
@@
Panels (c–e) shows that the estimate of $\Delta F$, using Eq. (13), becomes increasingly precise as $\lambda$ increases. The estimator $J_{\lambda}$ converges to $\beta\Delta F\approx 0$, and the error measures $\sigma^{2}_{B}$ and $N_{0.1}$ decrease by orders of magnitude. For fixed trajectory time $t_{\rm f}=1$, the $\lambda$-modified dynamics (using a simple linear protocol) allows us to estimate $\Delta F$ with much higher precision than does the original dynamics using the optimal protocol.
@@
In Fig. 1(f), the vertical dotted line denotes the value of $\beta\langle W\rangle_{\rm opt}\approx 8.33$, which is much larger than the mean value $\beta_{\lambda}\langle W_{\lambda}\rangle_{\lambda}\approx 0.8$ obtained using the $\lambda$-modified dynamics. For the latter, for $\lambda=30$, the estimate of $\beta\Delta F$ obtained using the Jarzynski equality is $J_{\lambda}\approx-0.000931229$, with error parameters $\sigma^{2}_{B}\approx 0.0394953$ and $N_{0.1}\approx 400$. Injecting noise has therefore increased the precision with which we can estimate $\Delta F$ using the Jarzynski equality.
@@
For the case $\lambda=1$, corresponding to the original dynamics, the protocol results in a far-from-equilibrium dynamics. For the case $\lambda=30$, the system remains close to equilibrium throughout the transformation. In the latter case the dissipated mean work $\beta_{\lambda}\langle W_{\lambda}\rangle_{\lambda}-\Delta F$ is smaller, and the resulting free-energy estimate more accurate. The snapshots illustrate the fact that adding noise to the system increases its relaxation rate, and scaling the potential accordingly leaves the thermodynamics of the system unchanged.

### v3

@@ Abstract
Estimating free-energy differences using nonequilibrium work relations, such as the Jarzynski equality, is hindered by poor convergence when work fluctuations are large. For systems governed by overdamped Langevin dynamics, we propose the counterintuitive approach of adding noise in order to increase the precision of such calculations. By introducing additional stochastic fluctuations to the system and rescaling its potential energy accordingly, we leave the thermodynamics of the system unchanged while increasing its relaxation rate. For a given time-dependent protocol this modification reduces the dissipated reduced work, leading to more accurate free-energy estimates. The method is designed to be used in experiment, and we illustrate its operation using computer simulations applied to two model systems. However, the regime of applicability of this strategy is likely limited, because it requires control of the system’s potential energy in a way that is feasible in only a few experimental settings.
@@ I Introduction
The Jarzynski equality and the Crooks fluctuation theorem provide a way to estimate equilibrium free-energy differences from nonequilibrium work measurements Jarzynski (1997); Crooks (1999). These relations apply to systems initially at equilibrium and driven out of it by the variation of a set of external control parameters. However, such estimates can be numerically challenging. The Jarzynski equality involves an exponential average, and rare trajectories with atypically small work values contribute disproportionately, making free-energy estimates unreliable unless a large number of trajectories are sampled Jarzynski (2006); Hummer and Szabo (2010); Rohwer et al. (2015). This problem worsens as the rate of driving increases, because the mean dissipated reduced work $\beta(\langle W\rangle-\Delta F)$ grows  [^1] The adjective reduced refers to the multiplication of a quantity by $\beta$., leading to an exponential increase in the number of trajectories required to estimate the free-energy difference to a given precision. Here $\beta\equiv 1/(k_{\rm B}T)$. Various strategies have been explored to mitigate this issue, including optimizing control protocols to minimize dissipated work and work fluctuations Schmiedl and Seifert (2007); Geiger and Dellago (2010); Vaikuntanathan and Jarzynski (2008); Blaber and Sivak (2020); Arrar et al. (2019); Loos et al. (2024); Zhong and DeWeese (2022).
@@
In this paper we propose an alternative and counterintuitive approach to increasing the precision of free-energy measurements for systems governed by overdamped Langevin dynamics: add noise. By injecting additional stochastic fluctuations into the system, and rescaling its potential energy accordingly, we can increase the system’s relaxation rate without altering its thermodynamics. For fixed driving rate, this modification reduces the reduced dissipated work and so suppresses rare-event sampling problems, leading to more precise estimates of $\Delta F$. We demonstrate this effect in two model systems, using numerical simulations, and show that this noise-engineering strategy can significantly enhance the accuracy of free-energy calculations. Our results indicate that noise engineering could be a valuable tool for enhancing thermodynamic measurements and computations, and suggest that the notion of optimal control Alvarado et al. (2025) could be broadened to include control of noise, which is now experimentally feasible Martinez et al. (2012); Chupeau et al. (2018); Saha et al. (2023); Aifer et al. (2024), as well as control of the potential.
@@
Our intent is that this method be applied to experiments, and we use numerical simulations to illustrate its operation (see Refs. Vaikuntanathan and Jarzynski (2008); Sivak et al. (2013); Wu et al. (2020); Zhong et al. (2024) for approaches designed to enhance sampling and precision in numerical settings). However, the regime of applicability of this method to experimental free-energy calculations is limited. Noise-injection methods have been demonstrated in various experimental settings, but the method discussed here also depends on our ability to control the system’s potential energy, which is feasible in only a few experimental settings. We describe the method in Section II, and apply it to two model systems in Section III. We conclude in Section IV, where we comment on the applicability of the approach to various experimental systems.
@@ II Free energies from nonequilibrium measurements
$\dot{x}_{i}=-\mu\frac{\partial}{\partial x_{i}}U({\bm{x}},{\bm{c}}(t))+\sqrt{2k_{\rm B}T\mu}\,\eta_{i}(t),$ (1)
on the time interval $t\in[0,t_{\rm f}]$. Here $\mu$ is the mobility parameter, which sets the basic time scale of the system, and $\eta_{i}$ is a Gaussian white noise with mean $\langle\eta_{i}(t)\rangle=0$ and covariance $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. The noise $\eta$ represents the thermal fluctuations inherent to the system.
@@
$P_{\rm F}(W)$ denotes the probability distribution of work under the forward protocol ${\bm{c}}(t)$; $P_{\rm R}(-W)$ denotes the probability distribution of the negative of the work under the time-reversed protocol ${\bm{c}}(t_{\rm f}-t)$ (starting in equilibrium with respect to the potential $U({\bm{x}},{\bm{c}}(t_{\rm f}))$); and the average $\langle\cdot\rangle$ is taken over many independent realizations of the forward protocol.
@@
where $i$ labels trajectories, as well as a measure of the statistical error in this quantity, the variance $\sigma^{2}_{B}$ of the block average $B$ of the Jarzynski estimator. The $j^{\rm th}$ block average
$B_{j}=N_{\rm block}^{-1}\sum_{i=1}^{N_{\rm block}}e^{-\beta W_{j,i}}$ (7)
is computed from $N_{\rm block}=100$ trajectories, where $W_{j,i}$ is the work associated with trajectory $i$ in block $j$. We estimate the variance $\sigma_{\rm B}^{2}$ of the block average as
$\sigma_{B}^{2}=N_{\rm samples}^{-1}\sum_{j=1}^{N_{\rm samples}}B_{j}^{2}-\overline{B}^{2},$ (8)
where
$\overline{B}=N_{\rm samples}^{-1}\sum_{j=1}^{N_{\rm samples}}B_{j},$ (9)
and $N_{\rm samples}=10^{4}$. The virtue of the block average is that it provides a straightforward way to estimate the statistical uncertainty of the Jarzynski estimator without requiring assumptions about the underlying distribution of $W$. By aggregating trajectories into independent blocks, we obtain a distribution of values $B$ from which the variance can be directly calculated.
We can also calculate the number of trajectories $N(\Delta f)$ required to estimate $\Delta F$ to a precision of $\Delta f\,k_{\rm B}T$, using the formula Geiger and Dellago (2010)
$N_{\Delta f}=\frac{1}{(\Delta f)^{2}}\left(\left\langle{\rm e}^{-2\beta W}\right\rangle-1\right).$ (10)
@@
$\dot{x}_{i}=-\mu\frac{\partial}{\partial x_{i}}\lambda U({\bm{x}},{\bm{c}}(t))+\sqrt{2k_{\rm B}T\mu}\,\eta(t)+\sigma\zeta_{i}(t),$ (11)
where $\langle\zeta_{i}(t)\rangle=0$ and $\langle\zeta_{i}(t)\zeta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. The two noise terms in (12) can be considered an effective Gaussian white noise of variance $2k_{\rm B}T\mu\lambda$, and so (11) can also be written
$\dot{x}_{i}=-\mu\frac{\partial}{\partial x_{i}}U_{\lambda}({\bm{x}},{\bm{c}}(t))+\sqrt{2k_{\rm B}T_{\lambda}\mu}\,\eta_{i}(t),$ (12)
@@
where $\langle\cdot\rangle_{\lambda}$ denotes an average over the $\lambda$-modified dynamics, Eq. (11). (In previous equations, the average $\langle\cdot\rangle$ corresponds to the original dynamics, i.e. to the choice $\lambda=1$.)
Next, note that the exponent appearing in the Jarzynski relation is not the work $W$ but the reduced work $\beta W$. Accordingly, the quantity in the exponential on the left-hand side of (14) is $\beta_{\lambda}W_{\lambda}$, appropriate for the scaled potential and effective temperature appearing in Eq. (12). The resulting estimate of $\beta\Delta F$ is
@@
This effective time-rescaled description suggests that the $\lambda$-modified dynamics will allow more accurate estimation of $\beta\Delta F$ using the Jarzynski identity than does the original dynamics. From our understanding of how work scales with trajectory duration Schmiedl and Seifert (2007); Van Vu and Hasegawa (2021) we expect the dissipated reduced work $\beta_{\lambda}\langle W_{\lambda}\rangle_{\lambda}-\beta\Delta F$ to decrease with $\lambda$, and for large $\lambda$ to scale as $1/\lambda$. Reducing the dissipated reduced work tends to reduce the variance of the Jarzynski estimator, and so we expect the statistics of $\langle{\rm e}^{-\beta_{\lambda}W_{\lambda}}\rangle_{\lambda}$ to become increasingly better behaved as $\lambda$ increases.
For example, for the illustrative case of Gaussian work fluctuations, discussed in Section A, a linear decrease with $\lambda$ in the dissipated reduced work results in a linear decrease with $\lambda$ in the variance of the dissipated reduced work. The latter results in an exponential decrease with $\lambda$ in the variance of the Jarzynski estimator, and hence an exponential decrease with $\lambda$ in the number of independent trajectories required to resolve the free-energy difference to a given precision.
@@ III.1 Particle in a trap
Figure 2: Trap-translation model using the linear protocol within the $\lambda$-modified dynamics (11) (green lines). The trajectory duration is $t_{\rm f}=1$, as in Fig. 1. Dashed lines labeled “opt” correspond to the optimal protocol used within the original dynamics (for which $\lambda=1$). (a) Mean work $\langle W_{\lambda}\rangle_{\lambda}$, in units of $k_{\rm B}T$, as a function of $\lambda$. (a) Mean reduced work $\beta_{\lambda}\langle W_{\lambda}\rangle_{\lambda}$ as a function of $\lambda$. As expected from the arguments connecting Eq. (11) and Eq. (16), this quantity behaves as $\beta\langle W\rangle$ for the original dynamics, Eq. (1), if the protocol is imposed $\lambda$ times more slowly for $\lambda$ times longer. The function $w$ is given by Eq. (19). (c) Free-energy estimator (15). (d) Variance of the block average (7), using $N_{\rm block}=100$ and $N_{\rm samples}=10^{4}$. (e) Number of trajectories (10) required to estimate $\Delta F$ to a precision $0.1k_{\rm B}T$; see Eq. (10). (f) Work distributions for forward and reverse protocols, for $\lambda=30$; compare the original dynamics of Fig. 1(b). Averages and distributions are calculated using $N_{\rm traj}=10^{6}$ trajectories.
@@
Figure 5: Erasure using the basic protocol of Ref. Dago et al. (2023) within the $\lambda$-modified dynamics (11). (a) Time-dependent protocol ${\bm{c}}(t)$. (b) Probability of successful erasure as a function of $\lambda$. (c) Free-energy estimator (15). (d) Variance of the block average (7). Averages were calculated using $10^{5}$ independent trajectories.
@@
For this system the work fluctuations are Gaussian, and so the work-minimizing protocol also minimizes work fluctuations and the error in estimating (3Geiger and Dellago (2010).
@@
which is slightly larger than that of the optimal protocol for all $t_{\rm f}$ Schmiedl and Seifert (2007). For large $t_{\rm f}$, both $\beta\langle W\rangle_{\rm opt}$ and $\beta\langle W\rangle_{\rm lin}$ go to zero as $\sim c_{\rm f}^{2}/t_{\rm f}$.
@@
The equation of motion for this system is linear stochastic differential equation with additive Gaussian white noise, making the trajectory $\{x(t)\}$ a Gaussian process. The total work performed during a protocol is a linear functional of this trajectory,
$W[x(t)]=-k\int_{0}^{t_{\rm f}}{\rm d}t\,(x(t)-c(t))\dot{c}(t),$ (20)
and, because linear functionals of Gaussian processes are Gaussian random variables, the work distribution $P(W)$ is Gaussian, regardless of $c(t)$. For Gaussian work fluctuations, as shown in Section A, the variance of the work is related to its mean by the expression
$\sigma^{2}=2k_{\rm B}T\langle W\rangle.$ (21)
For the linear protocol we therefore have $\sigma_{\rm lin}^{2}=2k_{\rm B}T\langle W\rangle_{\rm lin}$, with the mean given by Eq. (19).
@@
In Fig. 2 we show results obtained using the $\lambda$-modified dynamics (green lines), together with those obtained using the optimal protocol. Panel (a) shows that the mean work $\langle W_{\lambda}\rangle_{\lambda}$ increases with $\lambda$, as might be expected from the use of an energy function $\lambda$ times larger than the original. The new dynamics is more costly energetically than the original dynamics, even without accounting for the energy required to inject the added noise (which scales as $\sigma^{2}$ Whitelam (2024)).
However, consistent with our expectation from the previous section, the reduced work $\beta_{\lambda}\langle W_{\lambda}\rangle_{\lambda}$ decreases with increasing $\lambda$, because the system behaves as if it is subject to a slowed-down version of the original protocol. This decrease is shown in Fig. 2(b). Indeed, we see that
@@
Panels (c–e) show that the estimate of $\Delta F$, using Eq. (15), becomes increasingly precise as $\lambda$ increases. The estimator $J_{\lambda}$ converges to $\beta\Delta F\approx 0$, and the error measures $\sigma^{2}_{B}$ and $N_{0.1}$ decrease by orders of magnitude. For fixed trajectory time $t_{\rm f}=1$, the $\lambda$-modified dynamics (using a simple linear protocol) allows us to estimate $\Delta F$ with much higher precision than does the original dynamics using the optimal protocol.
@@
In Fig. 1(f), the vertical dotted line denotes the value of $\beta\langle W\rangle_{\rm opt}\approx 8.33$, which is much larger than the mean value $\beta_{\lambda}\langle W_{\lambda}\rangle_{\lambda}\approx 0.8$ obtained using the $\lambda$-modified dynamics. For the latter, for $\lambda=30$, the estimate of $\beta\Delta F$ obtained using the Jarzynski equality is $J_{\lambda}\approx-0.000931229$, with error parameters $\sigma^{2}_{B}\approx 0.0394953$ and $N_{0.1}\approx 400$. Injecting noise has therefore substantially increased the precision with which we can estimate $\Delta F$ using the Jarzynski equality.
@@
For the case $\lambda=1$, corresponding to the original dynamics, the protocol results in a far-from-equilibrium trajectory ensemble. For the case $\lambda=30$, the system remains close to equilibrium throughout the transformation. In the latter case the dissipated mean reduced work $\beta_{\lambda}\langle W_{\lambda}\rangle_{\lambda}-\beta\Delta F$ is smaller, and the resulting free-energy estimate more accurate. The snapshots illustrate the fact that adding noise to the system increases its relaxation rate, and scaling the potential accordingly leaves the thermodynamics of the system unchanged.